# An Auditable AI Agent Loop for Empirical Economics: A Case Study in Forecast Combination

### 19 Mar 2026 v2

@@ 2 An auditable agent-loop protocol
This distinction matters. $\tau^{\star}$ is the best admissible script in the full class $\mathcal{T}(C)$. $\hat{\tau}_{K}$ is the best script found by a particular bounded search. The gap between the two reflects the cost of finite search. The audit log records the outcome of every experiment as an ordered sequence $L_{K}=\bigl((k,\tau_{k},s_{k},d_{k})\bigr)_{k=0}^{K}$, where $s_{k}=S(\tau_{k};D^{S})$ when the run succeeds, $s_{k}=+\infty$ when it crashes, and $d_{k}\in\{\textit{keep},\,\textit{discard},\,\textit{crash}\}$ records the outcome status. Algorithm 1 summarizes the protocol.
@@ 3.1 Setup
We also evaluate the discovered methods on a post-search holdout (2017Q1–2025Q4, 36 quarters) that was withheld from the evaluator during the agentic search. As additional benchmarks, we include two look-ahead-free methods from DieboldShin2019: best average flex ($N_{\max}=6$, their Table 4) and best $(N,W)$ average adaptive (their Table 5), both of which exhaustively search over forecaster subsets at each forecast origin. Details on the holdout data construction are in Appendix E.
@@ 3.2 Search-sample and holdout results
Table 1 reports RMSE for the search sample (1999Q3–2016Q4) and the holdout (2017Q1–2025Q4). On the search sample, all three agent runs beat both the simple average and the peLASSO ex post benchmark: Run 1 achieves a relative RMSE of 0.858, Run 3 reaches 0.808, and Run 2 achieves 0.510. The two benchmark methods from the original paper also outperform the simple average, with relative RMSEs of 0.954 (best average flex) and 0.916 (best $(N,W)$ average adaptive).
@@
Best avg flex $1.435$ $0.954$ 2.901 [.240] $0.974$ 1.043 [.347] $0.932$
Best $(N,W)$ avg $1.378$ $0.916$ 2.901 [.127] $0.974$ 1.142 [.602] $1.020$
@@
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Best avg flex and best $(N,W)$ avg are from DieboldShin2019, Tables 4 and 5. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2026); these have intrinsically low power given the small holdout sample (36 quarters, 32 excluding COVID) and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post (original) uses $\lambda$ selected on the search sample for all windows; peLASSO ex post (per window) re-optimizes $\lambda$ on each evaluation window separately. Full results including intermediate methods are in Appendix E.
On the holdout, Run 2 remains the strongest method, with a relative RMSE of 0.811 (0.739 excluding the COVID quarters of 2020). Run 1 also holds up, at 0.945 (0.861 excluding COVID). The two benchmark methods from the original paper generalize modestly, each at 0.974. In contrast, Run 3 (1.089) and the peLASSO ex post original (1.065) perform worse than the simple average on the holdout; even the per-window variant that re-optimizes $\lambda$ on each holdout window achieves only 0.995, indicating that peLASSO’s difficulty on the holdout is not merely a lambda selection problem. The holdout therefore shows that agentic search can uncover methods that generalize, but it also reveals that some discovered improvements are sample specific.
@@ 3.3 What the log reveals
The instruction contract asks the agent to tune regularization parameters for peLASSO. The experiment logs [^2] Appendix C reproduces an excerpt from the experiment log of Run 3, and Appendix D gives a detailed description of the method discovered by each run., however, reveal that two of the three runs drifted beyond that task: Run 1 toward stability selection and performance weighting, and Run 2 toward a ranking-based method with bias correction. Run 3 stayed closest to the original task, using adaptive LASSO with forward cross validation for lambda selection, though it added an egalitarian elastic net blend. The holdout makes these departures evaluable. Run 2’s bias correction, which exploits serial correlation in combined forecast errors, turns out to generalize and accounts for most of its holdout gain. Run 3’s egalitarian elastic net blend, which drove much of its search-sample improvement, degrades holdout performance. The protocol does not prevent the agent from leaving the intended design space; it ensures that every departure is recorded, so that a holdout or other external evidence can later distinguish structural insight from sample adaptation.
@@ 4 Conclusion
Agent loops can widen researcher degrees of freedom, but they need not do so invisibly. This paper adapts a minimal agent loop architecture to an auditable protocol for agentic specification search and adds a post-search holdout evaluation, requiring only a written instruction contract, a fixed evaluator, a single editable script, a finite search budget, and a complete experiment log. In the forecast combination illustration based on DieboldShin2019, three independent runs improve on the search sample, but a genuine holdout shows that only two of the three discovered methods generalize. The contribution of the protocol is not to eliminate adaptive search, but to make it visible, bounded, and evaluable. Because the protocol disciplines the search process rather than the subject matter, it applies wherever automated specification search is used. Appendix F sketches how the same architecture maps to regression specification search, structural VAR identification, and other standard empirical tasks.
@@ D.2 Run 2: ranking, adaptive windows, and bias correction
    Select $N^{*}=\arg\min_{N}\sqrt{n^{-1}\sum_{i}(\hat{y}_{i}^{\mathrm{LOO}}-y_{i})^{2}}$.
@@
    Bias correction. The final forecast adjusts for serial correlation in forecast errors:
@@ D.3 Run 3: adaptive LASSO with forward CV and ensemble
    Multi horizon forward CV. For each fold $f$ with training end $t_{f}$, compute forecast errors at horizons $h=1$ and $h=2$. Two candidate aggregation rules are evaluated: simple average and blended average ($0.7\times\text{mean}+0.3\times\text{ridge weighted mean}$). The CV criterion is
    $\mathrm{CV}(\lambda)=\sum_{f}w_{f}\left(0.05\cdot e_{f,h=1}^{2}(\lambda)+0.95\cdot e_{f,h=2}^{2}(\lambda)\right),$ (9)
@@
Steps 1–2 as in Run 3. Forward CV with $h=1$ only and $\gamma=0.75$. No model averaging. Simple average of selected forecasters. No egalitarian elastic net blend:
@@
Steps 1–4 as in Run 3, except $\gamma=0.80$ and $\epsilon=10^{-3}$. Model averaging is applied. No egalitarian elastic net blend: $\hat{y}_{t}=\hat{y}_{t}^{\mathrm{PE}}$.
@@ E.1 Data and processing
E.2 Experiment counts
Table 3 reports the experiment counts for the three agent runs. Each run uses Claude Code with Opus 4.6 at default thinking effort, starting from a fresh agent with the same initial script (simple average) and an approximate budget of $K=200$ experiments. The evaluator calls the forecasting function 66 times ($t=5,\ldots,70$) but scores RMSE over 65 periods ($t=6,\ldots,70$); the first call serves as a warm-up. The budget is enforced as a soft constraint; the realized counts are 232, 228, and 201 for Runs 1, 2, and 3, respectively. If a candidate script exceeds 30 minutes of runtime, the experiment is terminated and logged as a crash.
Table 3: Experiment counts
Experiments Keeps Discards Crashes
Run 1 232 42 188 2
Run 2 228 61 164 3
Run 3 201 46 152 3
E.3 Full holdout results
@@
E.4 Role of individual components
@@
Removing the bias correction from Run 2 degrades performance from 0.811 to 0.990 (Run 2a) or 0.980 (Run 2b). The bias correction contributes most of Run 2’s advantage. However, Run 2a and 2b are still competitive with the simple average, indicating that the temporally weighted ranking and adaptive window mechanism also add modest value. The bias correction exploits a structural feature—serial correlation in combined forecast errors—rather than overfitting to the search sample.
@@
E.5 Independent dataset robustness
@@
Table 5 repeats the exercise on this independent dataset. The key rankings are preserved: Run 2 achieves a relative RMSE of 0.808 on the holdout, Run 1 achieves 0.946, and the original paper methods achieve 0.974 and 0.968. The only notable difference is Run 3, which improves from 1.089 to 0.962, suggesting its search-sample performance was sensitive to the specific GDP vintage. The independent dataset confirms that the main ranking—Run 2 strongest, Run 1 second—is preserved across vintages, while Run 3 is sensitive to the GDP vintage used.
@@
Appendix F Illustration: forecast combination and beyond
@@
  • program.md ($C$): “Compare candidate specifications of $g(z_{i})$ using a held-out evaluation sample. Only edit train.R. The evaluation sample partition is fixed in prepare.R. Do not modify the sample or the outcome variable.”
@@
  • prepare.R ($S$): Splits the data into a training sample $\mathcal{I}_{\text{train}}$ and an evaluation sample $\mathcal{I}_{\text{eval}}$, fixed ex ante. On $\mathcal{I}_{\text{train}}$, it sources train.R to obtain the nuisance fit, and if desired partials out both outcome and treatment before estimating $\hat{\beta}$. On $\mathcal{I}_{\text{eval}}$, it computes a holdout prediction error such as
    $S(\tau;D)=\frac{1}{|\mathcal{I}_{\text{eval}}|}\sum_{i\in\mathcal{I}_{\text{eval}}}\bigl(\log w_{i}-\hat{\beta}\cdot\text{educ}_{i}-\hat{g}(z_{i})\bigr)^{2}.$
  • results.tsv ($L_{K}$): Records each specification attempted, its holdout MSE, and the implied $\hat{\beta}$.
If the scoring rule were the in-sample $p$-value for $\hat{\beta}$, the loop would amount to automated $p$-hacking. The discipline comes from fixing the evaluator and the sample split before the search begins and from disclosing the full search path. Causal claims about returns to schooling would still require separate identification assumptions.
@@
  • program.md ($C$): “Compare candidate constructions of a monetary policy shock series using a fixed external diagnostic on a held-out evaluation sample. Only edit train.R. Do not modify the VAR specification or the evaluation sample.”
@@
  • prepare.R ($S$): Estimates the reduced-form VAR on a fixed training sample, calls train.R to obtain the candidate shock, and computes a diagnostic score on a held-out evaluation period:
    $S(\tau;D)=-\left|\text{corr}\!\left(\hat{\varepsilon}_{t}^{mp},\;z_{t}^{RR}\right)\right|,$
    where $z_{t}^{RR}$ is the Romer and Romer narrative monetary policy shock measure and the evaluation period is fixed ex ante. The negative sign ensures that higher correlation corresponds to a lower (better) score. This score is best read as an external diagnostic rather than a standalone proof of identification.
@@
The point is not that the agent “finds the largest effect” or settles identification, but that any search over candidate shock constructions is visible, bounded, and judged against a criterion fixed ex ante.
@@
    $S(\tau;D)=\text{RMSE}=\left(\frac{1}{|\mathcal{E}|}\sum_{t\in\mathcal{E}}\bigl(\pi_{t+h}-\hat{\pi}_{t+h|t}\bigr)^{2}\right)^{1/2}.$
@@
This is a natural fit for the protocol because the evaluator is a genuine pseudo-out-of-sample forecast loss. For central bank staff who routinely compare Phillips curve specifications, the protocol would make the search path more visible and easier to disclose.

### 20 Mar 2026 v3

@@ 2 An auditable agent-loop protocol
This distinction matters. $\tau^{\star}$ is the best admissible script in the full class $\mathcal{T}(C)$. $\hat{\tau}_{K}$ is the best script found by a particular bounded search. The audit log records the outcome of every experiment as an ordered sequence $L_{K}=\bigl((k,\tau_{k},s_{k},d_{k})\bigr)_{k=0}^{K}$, where $s_{k}=S(\tau_{k};D^{S})$ when the run succeeds, $s_{k}=+\infty$ when it crashes, and $d_{k}\in\{\textit{keep},\,\textit{discard},\,\textit{crash}\}$ records the outcome status. Algorithm 1 summarizes the protocol.
@@ 3.1 Setup
We also evaluate the discovered methods on a post-search holdout (2017Q1–2025Q4, 36 quarters) that was withheld from the evaluator during the agentic search. As additional benchmarks, we include two methods from DieboldShin2019: best $\leq N_{\max}$-average ($N_{\max}=6$, their Table 4) and best $(\leq N_{\max},\leq W_{\max})$-average (their Table 5), both of which exhaustively search over forecaster subsets at each forecast origin.
@@ 3.2 Search-sample and holdout results
Table 1 reports RMSE for the search sample (1999Q3–2016Q4) and the holdout (2017Q1–2025Q4). On the search sample, all three agent runs beat both the simple average and the peLASSO ex post benchmark: Run 1 achieves a relative RMSE of 0.858, Run 3 reaches 0.808, and Run 2 achieves 0.510. The two benchmark methods from the original paper also outperform the simple average, with relative RMSEs of 0.954 (best $\leq N_{\max}$-average) and 0.916 (best $(\leq N_{\max},\leq W_{\max})$-average).
@@
peLASSO ex post $1.400$ $0.930$ 2.964 [.191] $0.995$ 1.075 [.136] $0.960$
Best $\leq 6$-avg $1.435$ $0.954$ 2.901 [.240] $0.974$ 1.043 [.347] $0.932$
Best $(\leq 6,\leq 40)$-avg $1.378$ $0.916$ 2.901 [.127] $0.974$ 1.142 [.602] $1.020$
@@
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Best $\leq 6$-avg and best $(\leq 6,\leq 40)$-avg are from DieboldShin2019, Tables 4 and 5. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2026); these have intrinsically low power given the small holdout sample and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post re-optimizes $\lambda$ on each evaluation window separately; see Appendix E for the fixed-$\lambda$ variant. Full results including intermediate methods are in Appendix E.
On the holdout, Run 2 remains the strongest method, with a relative RMSE of 0.811 (0.739 excluding the COVID quarters of 2020). Run 1 also holds up, at 0.945 (0.861 excluding COVID). The two benchmark methods from the original paper generalize modestly, each at 0.974. In contrast, Run 3 (1.089) performs worse than the simple average on the holdout, and peLASSO ex post shows essentially no improvement (0.995). The holdout therefore shows that agentic search can uncover methods that generalize, but it also reveals that some discovered improvements are sample specific.
@@ 3.3 What the log reveals
The instruction contract asks the agent to tune regularization parameters for peLASSO. The experiment logs [^2] Appendix C reproduces an excerpt from the experiment log of Run 3, and Appendix D gives a detailed description of the method discovered by each run., however, reveal that two of the three runs drifted beyond that task: Run 1 toward stability selection and performance weighting, and Run 2 toward a ranking-based method with bias correction. Run 3 stayed closest to the original task, using adaptive LASSO with forward cross validation for lambda selection, though it added an egalitarian elastic net blend. The holdout makes these departures evaluable. Run 2’s bias correction, which adjusts for recent forecast bias, turns out to generalize and accounts for most of its holdout gain. Run 3’s egalitarian elastic net blend, which drove much of its search-sample improvement, degrades holdout performance. The protocol does not prevent the agent from leaving the intended design space; it ensures that every departure is recorded, so that a holdout or other external evidence can later distinguish structural insight from sample adaptation.
@@ 4 Conclusion
Agent loops can widen researcher degrees of freedom, but they need not do so invisibly. This paper adapts a minimal agent loop architecture to an auditable protocol for agentic specification search and adds a post-search holdout evaluation. In the forecast combination illustration based on DieboldShin2019, three independent runs improve on the search sample, but a genuine holdout shows that only two of the three discovered methods generalize. The contribution of the framework is not to eliminate adaptive search, but to make it visible, bounded, and evaluable. Because the protocol disciplines the search process rather than the subject matter, it applies wherever automated specification search is used. Appendix F sketches how the same architecture maps to regression specification search, structural VAR identification, and other standard empirical tasks.
@@ A.3 Isolating the holdout
A.4 Actual run
Each run uses Claude Code with Opus 4.6 at default thinking effort, starting from a fresh agent with the same initial script (simple average) and an approximate budget of $K=200$ experiments. The evaluator calls the forecasting function 66 times ($t=5,\ldots,70$) but scores RMSE over 65 periods ($t=6,\ldots,70$), matching the burn-in and evaluation design in DieboldShin2019. Table 2 reports the experiment counts for the three agent runs. The budget is enforced as a soft constraint; the realized counts are 232, 228, and 201 for Runs 1, 2, and 3, respectively. If a candidate script exceeds 30 minutes of runtime, the experiment is terminated and logged as a crash.
Table 2: Experiment counts
Experiments Keeps Discards Crashes
Run 1 232 42 188 2
Run 2 228 61 164 3
Run 3 201 46 152 3
@@ D.1 Run 1: stability selection with performance weighting
    The exponent $p=14$ was selected by the search from nearby alternatives (powers 2 through 20 were tried).
@@ D.2 Run 2: ranking, adaptive windows, and bias correction
    Select $N^{*}=\arg\min_{N}\sqrt{n^{-1}\sum_{i}(\hat{y}_{i}^{\mathrm{LOO}}-y_{i})^{2}}$. After $N^{*}$ is selected, the current period forecast (step 5) uses the full window weighted MAE ranking, not the LOO reranking or the $\mathrm{MAE}^{-3}$ weighted average.
@@
    Bias correction. The final forecast applies a bias adjustment based on the most recent leave one out error:
@@ D.3 Run 3: adaptive LASSO with forward CV and ensemble
    Multi horizon forward CV. Lambda is selected by walk forward CV on the full history preceding the current origin, rows $1,\ldots,t-1$. For fold $f$, fit adaptive LASSO on the first $t_{f}$ rows and compute forecast errors on the next two rows $t_{f}+1$ and $t_{f}+2$ (where $t_{f}+2\leq t-1$ by construction). Let $e_{f,s}(\lambda)=\hat{y}_{t_{f}+s}(\lambda)-y_{t_{f}+s}$ for $s=1,2$, where $\hat{y}_{t_{f}+s}(\lambda)$ applies the $\lambda$ dependent combination rule to the forecaster predictions at row $t_{f}+s$. Two candidate aggregation rules are evaluated at each $\lambda$: simple average and blended average ($0.7\times\text{mean}+0.3\times\text{ridge weighted mean}$). The CV criterion is
    $\mathrm{CV}(\lambda)=\sum_{f}w_{f}\left(0.05\cdot e_{f,1}^{2}(\lambda)+0.95\cdot e_{f,2}^{2}(\lambda)\right),$ (9)
@@
Steps 1–2 as in Run 3. Forward CV evaluating only at the next panel row ($t_{f}+1$) with $\gamma=0.75$. No model averaging. Simple average of selected forecasters. No egalitarian elastic net blend:
@@
Same pilot and adaptive penalty construction as Run 3, with $\gamma=0.80$ and $\epsilon=10^{-3}$. Forward CV evaluates only the simple average aggregation rule (no blended alternative). Model averaging over $\lambda$ values within 1% of the best CV score is applied. No egalitarian elastic net blend: $\hat{y}_{t}=\hat{y}_{t}^{\mathrm{PE}}$.
@@ E.1 Data and processing
E.2 Full holdout results
@@
E.3 Role of individual components
@@
Removing the bias correction from Run 2 degrades performance from 0.811 to 0.990 (Run 2a) or 0.980 (Run 2b). The bias correction contributes most of Run 2’s advantage. However, Run 2a and 2b are still competitive with the simple average, indicating that the temporally weighted ranking and adaptive window mechanism also add modest value. The bias correction adjusts for persistent short run bias in combined forecast errors, which appears to reflect a structural feature rather than overfitting to the search sample.
@@
E.4 Independent dataset robustness
@@
Table 5 repeats the exercise on this independent dataset. The key rankings are preserved: Run 2 achieves a relative RMSE of 0.808 on the holdout, Run 1 achieves 0.946, and the original paper methods achieve 0.974 and 0.968. Run 3’s relative RMSE shifts from 1.089 to 0.962 on the independent dataset, but its $p$-values remain very large in both cases, so the difference is not meaningful. The independent dataset confirms that the main ranking is preserved across vintages: Run 2 strongest, Run 1 second, and Run 3 not distinguishable from the simple average.
@@
Appendix F Illustration: beyond forecasting combination
@@
  • program.md ($C$): “Compare candidate specifications of $g(z_{i})$. Only edit train.R. The search sample and its internal evaluation split are fixed in prepare.R. Do not modify the sample or the outcome variable.”
@@
  • prepare.R ($S$): Splits the search sample $D^{S}$ into a training partition $\mathcal{I}_{\text{train}}$ and an evaluation partition $\mathcal{I}_{\text{eval}}$, fixed ex ante. On $\mathcal{I}_{\text{train}}$, it sources train.R to obtain the nuisance fit, and if desired partials out both outcome and treatment before estimating $\hat{\beta}$. On $\mathcal{I}_{\text{eval}}$, it computes a holdout prediction error such as
    $S(\tau;D^{S})=\frac{1}{|\mathcal{I}_{\text{eval}}|}\sum_{i\in\mathcal{I}_{\text{eval}}}\bigl(\log w_{i}-\hat{\beta}\cdot\text{educ}_{i}-\hat{g}(z_{i})\bigr)^{2}.$
  • results.tsv ($L_{K}$): Records each specification attempted, its evaluation MSE, and the implied $\hat{\beta}$.
If the scoring rule were the in-sample $p$-value for $\hat{\beta}$, the loop would amount to automated $p$-hacking. The discipline comes from fixing the evaluator and the sample split before the search begins and from disclosing the full search path. Causal claims about returns to schooling would still require separate identification assumptions. After the search, the researcher can evaluate the winning specification on a post-search holdout $D^{H}$ (e.g., a later survey year or a pre-reserved subsample withheld from the search) to check whether the nuisance fit generalizes.
@@
  • program.md ($C$): “Compare candidate constructions of a monetary policy shock series using a fixed external narrative diagnostic. Only edit train.R. Do not modify the VAR specification.”
@@
  • prepare.R ($S$): Estimates the reduced-form VAR on a fixed training sample, calls train.R to obtain the candidate shock, and computes a diagnostic score on $D^{S}$:
    $S(\tau;D^{S})=-\left|\text{corr}\!\left(\hat{\varepsilon}_{t}^{mp},\;z_{t}^{RR}\right)\right|,$
    where $z_{t}^{RR}$ is the Romer and Romer narrative monetary policy shock measure. The negative sign ensures that higher correlation corresponds to a lower (better) score. This score is best read as an external diagnostic rather than a standalone proof of identification.
@@
The point is not that the agent “finds the largest effect” or settles identification, but that any search over candidate shock constructions is visible, bounded, and judged against a criterion fixed ex ante. After the search, the researcher can evaluate the chosen shock construction on a later sample period $D^{H}$ (e.g., search on 1969–2000, holdout on 2001–2007) to check whether the narrative correlation and impulse response patterns hold outside the search sample.
@@
    $S(\tau;D^{S})=\text{RMSE}=\left(\frac{1}{|\mathcal{E}|}\sum_{t\in\mathcal{E}}\bigl(\pi_{t+h}-\hat{\pi}_{t+h|t}\bigr)^{2}\right)^{1/2}.$
@@
This is a natural fit for the protocol because the evaluator is a genuine pseudo-out-of-sample forecast loss. For central bank staff who routinely compare Phillips curve specifications, the protocol would make the search path more visible and easier to disclose. The researcher can further extend the evaluation to a later period $D^{H}$ not available during the search to check whether the selected specification generalizes.